OpenStudy (anonymous):

How would you go about establishing the identity of (tan u)(cot u)-cos^2u=sin^2u.

OpenStudy (jdoe0001):

do you know what ahemm say \(\bf sin^2(\theta)+cos^2(\theta)=?\)

OpenStudy (anonymous):

Yes, the Pythagorean identity that equals one.

OpenStudy (amistre64):

and what does tan*cot equal?

OpenStudy (anonymous):

1?

OpenStudy (jdoe0001):

well.... yes is 1 so \(\bf sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)={\color{brown}{ 1-cos^2(\theta)}} \\ \quad \\ cot(\theta)={\color{purple}{ \cfrac{1}{tan(\theta)}}} \\ \quad \\ \quad \\ tan(u)cot(u)-cos^2(u)=sin^2(u)\\ \quad \\ \cancel{ tan(u) }\cdot {\color{purple}{ \cfrac{1}{\cancel{ tan(u) }}}}-cos^2(u)=sin^2(u) \\ \quad \\ {\color{brown}{ 1-cos^2(u)}}=sin^2(u)\)

OpenStudy (anonymous):

Okay, that makes sense. I worked it out in a somewhat similar way, but I messed up on the step where you crossed out the tan u. Thank you so much for showing me how to find the other side.